OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2+6x-7=0. @Taylor31766

OpenStudy (anonymous):

Pls i need help with this question. thanks

OpenStudy (anonymous):

x^2 - 6x - 7 = 0 (x-7)(x+1) So x = 7; x = -1 7^2 - (6)(7) - 7 = 0 49 - 42 - 7 = 0 0 =0 (-1)^2 - (6)(-1) - 7 1 + 6 - 7 = 0 0 = 0 So, they check.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

npnp

OpenStudy (anonymous):

Solve the quadratic equation by completing the square. -x^2+x-8=0. thanks @Taylor31766

OpenStudy (anonymous):

Can't be bothered going another. Find it yourself?

OpenStudy (anonymous):

Ok